OpenStudy (anonymous):

Which property would allow you to use mental computation to simplify the problem 27 + 15 + 3 + 5? a,commutative property of addition b,distributive property c,additive identity d,additive inverse please help me:(

jimthompson5910 (jim_thompson5910):

Hint: 27 + 15 + 3 + 5 may be easier to do in your head if you swap 15 and 3 to get 27 + 3 + 15 + 5

OpenStudy (precal):

commutative propery commutative property of addition a+b=b+a commutative property of multiplication ab=ba

OpenStudy (precal):

using words commutative property allows you to add numbers in any order a+b=b+a or 12+3=3+12 for the second case it allows you to multiply numbers in any order ab=ba 2*5=5*2

OpenStudy (precal):

jim_thompson5910 is correct

OpenStudy (anonymous):

so it b???

jimthompson5910 (jim_thompson5910):

re-read what precal wrote

OpenStudy (anonymous):

ow its c

jimthompson5910 (jim_thompson5910):

you sure?

OpenStudy (anonymous):

no not really:(

jimthompson5910 (jim_thompson5910):

which property did precal mention?

OpenStudy (anonymous):

comunitive

jimthompson5910 (jim_thompson5910):

that's your answer